### v1

# A Survey on Human-Centered Evaluation of Explainable AI Methods in Clinical Decision Support Systems

Abstract
Explainable AI (XAI) has become a crucial component of Clinical Decision Support Systems (CDSS) to enhance transparency, trust, and clinical adoption. However, while many XAI methods have been proposed, their effectiveness in real-world medical settings remains underexplored. This paper provides a survey of human-centered evaluations of Explainable AI methods in Clinical Decision Support Systems. By categorizing existing works based on XAI methodologies, evaluation frameworks, and clinical adoption challenges, we offer a structured understanding of the landscape. Our findings reveal key challenges in the integration of XAI into healthcare workflows and propose a structured framework to align the evaluation methods of XAI with the clinical needs of stakeholders.
1 Introduction
Artificial Intelligence (AI) is increasingly being integrated into Clinical Decision Support Systems (CDSS) to improve timely diagnosis, optimize treatment plans, and reduce healthcare costs Sendak et al. (2020). Estimates suggest that AI adoption in healthcare could reduce U.S. healthcare expenditures by 5-10%, translating to \$200–\$360 billion in annual savings while maintaining quality and access Sahni et al. (2023). The rapid growth of AI-driven CDSS has also attracted significant investment, with venture capital funding in AI in health exceeding \$11 billion in 2024 Goldsack and Overgaard (2024). These trends indicate that the role of AI in medical decision making will continue to expand.
However, a key challenge accompanying this expansion is the necessity for trust and transparency in AI-driven medical decisions. Despite advancements in deep learning and other black-box AI models, clinicians often hesitate to adopt AI-generated recommendations due to their lack of interpretability Hou et al. (2024). Trust in CDSS is critical because errors in automated decision-making can have life-threatening consequences Zhang and Zhang (2023). In response, Explainable AI (XAI) has emerged as a solution to improve model transparency, inderectly allowing medical professionals to better understand why an AI system makes certain predictions Loh et al. (2022). Additionally, regulatory frameworks, such as the General Data Protection Regulation (GDPR), emphasize the need for explainability in automated decision-making, thereby necessitating interpretable AI solutions in healthcare.
Despite extensive research in XAI for healthcare Loh et al. (2022), a critical gap remains: there is limited systematic analysis of how XAI explanations impact real-world clinical adoption and decision-making Ghassemi et al. (2021). While prior work has primarily focused on developing new XAI techniques, fewer studies have systematically examined whether these methods effectively support clinicians in practice Amann et al. (2022). Furthermore, the evaluation of XAI in healthcare is highly fragmented, relying on inconsistent methodologies that range from proxy simulations to real-world clinical trials Doshi-Velez and Kim (2017). This raises a fundamental question: How should XAI techniques be evaluated to ensure their usability, trustworthiness, and effectiveness in real clinical settings?
To address this gap, this paper presents a systematic survey of human-centered evaluations of XAI methods in CDSS, focusing on the intersection of explainability, usability, and clinical decision-making. Specifically, we: (1) categorize existing works based on XAI methodologies, evaluation frameworks, and clinical adoption challenges, offering a structured synthesis of the field; (2) identify key gaps in current evaluation practices from clinicians perceptions of XAI-based CDSS, revealing that while XAI is generally well-received, significant challenges remain—such as high cognitive load, misalignment with clinical knowledge, and lack of early involvement with stakeholders; (3) introduce a conceptual framework to bridge the socio-technical gap in AI-driven CDSS by aligning XAI evaluation methods with the needs of diverse stakeholders (e.g., clinicians, hospital administrators, AI developers); (4) propose a medical field-agnostic, stakeholder-centric approach for CDSS development, advocating that CDSS should be designed as human-augmentation tools rather than replacements.
To theoretically ground this misalignment between human expectations and technological capabilities, we leverage Ackerman’s socio-technical gap theory, which highlights the challenges arising from such discrepancies Ackerman (2000). This analysis informs our proposed framework, emphasizing stakeholder-driven evaluations to ensure CDSS solutions are clinically relevant, interpretable, and effectively integrated into healthcare workflows. By aligning technical advancements in XAI with rigorous human-centered evaluations, we aim to steer research toward developing trustworthy, deployable CDSS solutions.
We organize the paper as follows. Section 2 provides background on CDSS, XAI methods, and human-centered evaluations. Section 3 presents a structured review of existing works in this field, categorizing them based on XAI methodologies, evaluation frameworks, and clinical adoption challenges. Section 4 discusses the socio-technical challenges and proposes a conceptual framework to bridge the gap between XAI methodologies and stakeholder needs, concluding pointing out future research directions.
2 Background
2.1 Clinical Decision Support Systems (CDSS)
CDSS have long been integral to medical decision-making, assisting clinicians in enhancing patient outcomes. Initially emerging in the late 1950s with rule-based systems, CDSS evolved significantly with advancements in AI. Modern CDSS now leverage sophisticated machine learning algorithms to process vast datasets efficiently. Concurrently, the rise of XAI has enabled the integration of interpretable solutions, essential for fostering trust and promoting the adoption of CDSS in clinical practice Ghassemi et al. (2021).
Operationally, CDSS utilize patient data to generate specific treatment recommendations Musen et al. (2021). The back-end employs machine learning algorithms trained for specific tasks, processing data from Electronic Health Records (EHRs). EHRs include both structured data (e.g., tabular data) and/or unstructured data (e.g., text, images), representing comprehensive digital versions of patients’ medical histories, including diagnoses, treatments, and test results. System performance is evaluated using machine learning metrics pertinent to the specific task. Synchronously, the front-end runs a user interface (UI) connected to the back-end that clinical practitioners are supposed to interact with. Such UI displays patient data, alerts, recommendations (sometimes with explanations), and decision-support reports in an accessible manner to facilitate clinical decision-making. To this end, human-centered evaluations are conducted to assess the overall effectiveness and efficiency of the system in the specific task. However, researchers have identified that the absence of such evaluations is a primary factor in the lack of adoption of AI-based CDSS solutions  Musen et al. (2021).
In summary, the integration of XAI together with human-centered evaluation assures that the system is usable within clinical workflows and more transparently meets the needs of stakeholders to foster acceptance of CDSS.
2.2 Explainability and XAI Methods: a Taxonomy
From a psychological standpoint, humans seek explanations to predict future events, as explanations facilitate generalization Vasil and Lombrozo (2022). Unlike descriptions, explanations provide understanding by identifying “difference-makers” in causal relationships. In medicine, where transparency and accountability are crucial, explanations may offer insights into the rationale behind AI-based CDSS recommendations, enabling practitioners to align the model’s reasoning with their medical expertise.
In AI, explainability and interpretability improve people’s understanding of such models, driving from suboptimal to “better” outcomes Boyd-Graber et al. (2022). More in general, the main principle of XAI is to make the decision-making process of AI systems transparent and understandable to humans. Many taxonomies have been proposed to categorize XAI methodologies (e.g., see Speith (2022)). In this paper, we adopt the conceptual approach based on the following macro criteria (for a more comprehensive review, please refer to Molnar (2022)):
  • Type: Intrinsic (or Ante Hoc) or Post Hoc;
  • Dependency: Model Specific or Model Agnostic;
  • Scope: Local or Global;
  • Output: Numerical, or Visual.
2.2.1 Type: Intrinsic or Post Hoc
Intrinsic interpretability is associated with machine learning models such that humans can easily trace their decision-making process. This is typically achieved through common white-box models such as linear models, in which coefficients can be considered the marginal effects on the output in regression problems, and tree methods as decision trees, in which their hierarchical structure allows for direct observation and understanding of the decision-making process. In addition, ensemble models such as random forests or boosting methods, although considered as black box, contain an intrinsic interpretable component in terms of feature contribution. Finally, some deep learning models may have intrinsic interpretability components in their structure, e.g., attention weights in Transformers, highlighting the importance of different input features. Methods: Linear Models, Tree-models, Ensemble Models, Attention Weights in Transformers, Bayesian Networks, etc.
Post Hoc interpretability refers to methodologies used after model training, providing insights into how predictions are made without altering the model itself. These methodologies are not intrinsic to the model, requiring further optimization to extract interpretation-oriented representations such as feature importance scores, partial dependence plots, surrogate models, among others. Notwithstanding, these methods can be applied on top of intrinsically interpretable models to provide a further layer of explainability as well. Methods: SHAP Lundberg and Lee (2017), LIME Ribeiro et al. (2016), GradCAM Selvaraju et al. (2017), DeepTaylor Montavon et al. (2017), Integrated Gradients Sundararajan et al. (2017), Surrogate Models, Partial Dependence Plots, Counterfactual Explanations, etc.
2.2.2 Dependency: Model Specific or Model Agnostic
Model Specific methods are limited to certain families of model architectures, and their use is coupled to the inner workings of those models. For example, neural networks require specific explainability methodologies, e.g., saliency maps such as GradCAM for CNNs. Importantly, note that intrinsic interpretable models are by definition model specific, e.g., interpreting linear regression coefficients is specific to that kind of model only. Methods: Architecture-specific Saliency Maps, GradCAM, DeepTaylor, Linear Models, Tree Paths, etc.
Model Agnostic methodologies are suitable for multiple families of machine-learning models, regardless of their internal structure. These methods are called “agnostic” because they do not rely on the internal workings of the model, which they treat as black box, but focus on the inputs and outputs only. These methodologies are applied post-hoc. Methods: SHAP, LIME, Partial Dependence Plots, Accumulated Local Effects, Counterfactual Explanations, Scoped Rules, Integrated Gradients, etc.
2.2.3 Scope: Local or Global
Local interpretability methods explain an individual prediction, e.g., how each feature contributes to a given output, thus approximating model’s behavior locally. Also, local methods are particularly suitable for models leveraging unstructured data, as, for instance, they can pinpoint the specific pixels in an image or tokens in a text that predominantly drive predictions. Methods: Individual Conditional Expectations, LIME, SHAP, Counterfactual Explanations, Integrated Gradients, Attention Weights, etc.
Global interpretability methods provide a wider perspective on how a model is explainable across the whole dataset (usually the test set). These methods are particularly suitable for models leveraging structured data. While local interpretability is essential for justifying individual choices, global interpretability promotes global model transparency. Methods: SHAP, Partial Dependence Plots, Accumulated Local Effects, Global Surrogates, etc.
2.2.4 Output: Numerical or Visual
Numerical explanations are suitable for structured data, especially for feature-based approaches. Examples include feature summary statistics, which is particularly indicated for intrinsic interpretable models such as linear regression, where coefficients, standard errors, and t-statistics provide information on how a particular feature is contributing to the model output, or tree-based methods such as decision trees, where feature importances can be traced back from the top of the tree. Besides, post hoc and model-agnostic methods such as SHAP or LIME can provide a feature-wise numerical value for pattern interpretations. Methods: all feature summary statistics methodologies developed for models leveraging structured data.
Visual output methods are particularly suitable for deep learning models operating with unstructured data such as images. Visualization techniques, such as saliency maps, class activation maps (CAM), and Grad-CAM, enable the interpretation of feature importance in CNNs, highlighting regions of interest in input images. Parallely, this also extends to token importance visualisation for language models. Beyond deep learning, structured data can also benefit from visual representations, including partial dependence plots, feature importance bar charts, and decision boundary plots. However, we treat those methodologies as numerical. Methods: all methodologies suitable for unstructured data.
2.3 Human-Centered XAI Evaluations
In practical medical applications, socio-technical gaps may arise between the CDSS explainability factors provided by XAI techniques and end-users’ perceptions of their utility Ackerman (2000). Human-centered evaluations offer methodologies to bridge this gap, aiming to align explanations with user expectations. Such evaluations should hopefully reflect human desired properties of explanations, serving some practical end goal Liao and Xiao (2023).
Human-centered evaluations of XAI systems can be categorized into three levels: (1) application-grounded, (2) human-grounded, and (3) proxy evaluation  Doshi-Velez and Kim (2017). Application-grounded evaluations involve testing real tasks with real experts and should be employed whenever possible because of their fidelity to real-world scenarios. For example, in the medical domain, doctors may be evaluated on specific application tasks. Here, domain experts are required for an effective evaluation. For this reason, application-grounded evaluation may be costly to implement. On the other hand, human-grounded evaluations, use proxy tasks with human evaluations, conducting simplified human-subject experiments. This approach is particularly suitable for assessing general aspects of a system. As such evaluations can be performed with layman participants instead of domain experts, they are not advisable in the medical domain. Finally, proxy evaluations assess proxy tasks without human participation, relying on simulated tasks to provide feedback about a system.
The main difference between application-grounded and human-grounded evaluation is the non-reliance on human subjects. Similar to human-grounded evaluations, proxy evaluations are not advisable for the full evaluation of CDSS ready for production. However, they may be useful for an early-stage prototype, provided a careful application-grounded evaluation is eventually performed. These three levels of XAI evaluations imply a trade-off between fidelity and cost, where application-grounded evaluations provide the most realistic insights but are expensive, human-grounded evaluations offer a balance of generalizability and affordability but lack domain specificity, and proxy evaluations are the most scalable and cost-efficient but least reflective of real-world performance. In the medical domain, application-grounded evaluation is the preferred choice because it ensures high-fidelity testing with domain experts. This is crucial for patient safety and clinical reliability, despite its higher cost and complexity.
Human-centered evaluation methodologies leverage various techniques, either qualitative or quantitative. In a nutshell, the most widely known are: (1) think-aloud studies, (2) interviews, and (3) surveys. First, in think-aloud studies participants verbalize their thoughts while performing tasks (concurrent) or afterward (retrospective), which can be very useful in qualitatively explaining cognitive processes and identifying usability issues. Second, interviews are a qualitative way to gather detailed information from users about their experiences with a system. They are versatile tools in all stages of design, ranging from open-ended unstructured interviews to tightly controlled structured interviews, with semi-structured interviews offering a valuable compromise. A variant of those is focus groups, in which a small number of participants are gathered to evaluate their opinions about a product or service. Finally, survey methodologies mainly rely on quantitative questionnaires that help researchers gather data on users’ satisfaction, usability issues, and interaction patterns.
3 Dissecting extant Human-Centered XAI in CDSS
In this section, we present research works that leveraged human-centered XAI to build production-ready CDSS in the medical field. Leveraging the XAI taxonomy outlined before, we discuss papers, highlighting how human-centered evaluations - all relying on application-grounded evaluations - were useful in testing clinical perceptions. In table 1, we provide a summary of the papers discussed. We start discussing CDSS adopting intrinsic interpretable models in Section 3.1. Next, we examine CDSS using post hoc and model agnostic methodologies on top of intrinsic interpretable models in Section 3.2. Finally, we discuss CDSS adopting XAI methodologies in deep learning in Section 3.3, starting with model-agnostic methods, and concluding with model-specific ones.
Paper Model XAI Type Dependency Output Main HCE
Barda et al. (2020) Random Forest SHAP Post Hoc Agnostic Numerical Focus Group
Brennan et al. (2019) Random Forest - Intrinsic Specific Numerical Think Aloud
Ellenrieder et al. (2023) DenseNet IGrads Post hoc Agnostic Visual Think Aloud
Hwang et al. (2022) TRIER (CNN) SG Post Hoc Specific Visual Mixed
Kim et al. (2022) XGBoost SHAP Post Hoc Agnostic Numerical Interviews
Kumarakulasinghe et al. (2020) Random Forest LIME Post Hoc Agnostic Numerical Survey
Kovalchuk et al. (2022) Many SHAP Post Hoc Agnostic Numerical Mixed
Kyrimi et al. (2020) Bayesian Network - Intrinsic Specific Numerical Survey
Matthiesen et al. (2021) Random Forest LIME Post Hoc Agnostic Visual Interviews
Neves et al. (2021) Many Many Post Hoc Agnostic Visual Survey
Pumplun et al. (2023) DenseNet IGrads Post Hoc Agnostic Visual Survey
Rajashekar et al. (2024) Random Forest ICE, ALE Post Hoc Agnostic Numerical Mixed
Sabol et al. (2020) CNN X-CFCMC Post Hoc Agnostic Visual Survey
Singh et al. (2021) Inception-v-3 DeepTaylor Post Hoc Specific Visual Survey
Sivaraman et al. (2023) XGBoost SHAP Post Hoc Agnostic Numerical Think Aloud
Zhang et al. (2024) LSTM Attention Intrinsic Specific Numerical Interviews
Table 1: Summary of papers in alphabetical order adopting XAI methods and Human-Centered Evaluations in medical CDSS. Categorization is executed by defining ML models, XAI methods and their taxonomy (type, dependency, output), and HCE methodologies.
3.1 CDSS adopting Intrinsic Interpretable Models
Intrinsic interpretable models are easily testable in the medical field for CDSS due to their transparency, as no post-hoc evaluations are necessarily required to perform model evaluations. Examples in this scope are the works from Brennan et al. (2019) and Kyrimi et al. (2020). Brennan et al. (2019) validated the usability of MyRiskSurgery, an analytical CDSS based on generalized linear models and random forests to estimate postoperative risk complications. Through think-aloud studies, a set of physicians evaluated risk complications for a set of 8-10 individual cases in two case studies such as diagnosing acute kidney injury and intensive care unit admission, with the system displaying the top 3 contributing features for each prediction. As a result, physicians experienced notable improvements in their risk assessment, which led to a substantial net improvement in reclassification rates from a no XAI evaluation (+12% and +16%, respectively), also finding the system user-friendly and beneficial.
Then, Kyrimi et al. (2020) examined how explanations from Bayesian Networks could be used to predict coagulopathy in the first 10 minutes of hospital care. Bayesian networks generate explanations by using probabilistic reasoning to identify the most likely causes based on the relationships between variables. In their framework, the tool outputted explanations in three levels of complexity ranking the predictors that supported a high risk of coagulopathy. In a before-after survey questionnaire, a group of clinicians rated the perception of explanations of the system, resulting in clinicians overall trusting the CDSS prediction, yet their level of trust remained unchanged when an explanation was provided, with the third level of explanations deemed too complex.
3.2 CDSS adopting Post Hoc XAI Methodologies on top of Intrinsic Interpretable Models
Post-hoc methodologies can be applied on top of intrinsic interpretable models to provide a further layer of explainability. In this subfield, we surveyed the works of Barda et al. (2020), Kim et al. (2022), Kumarakulasinghe et al. (2020), Kovalchuk et al. (2022), Matthiesen et al. (2021), Rajashekar et al. (2024), and Sivaraman et al. (2023). All these works employed model-agnostic methodologies.
Barda et al. (2020) developed a CDSS to prevent in-hospital mortality in pediatric intensive care units (PICU). The CDSS leveraged a random forest, and SHAP values on top, to predict PICU admission. Overall, feedback from focus group sessions with clinicians indicated that the proposed solution was perceived as useful. However, the capacity to digest the information varied across clinical roles and levels of predictive modeling knowledge. In general, they preferred solutions that required less cognitive effort and could accommodate diverse user information needs. Still adopting ensemble methods, Kumarakulasinghe et al. (2020) analyzed the agreement rate between physicians and LIME feature importances for sepsis prediction. Firstly, the authors trained several models such as random forest, adaboost, among others, and applied the LIME framework on top of the best-performing one (random forest). Secondly, physicians were asked to: (1) agree with the model whether the patient presented sepsis, based on individual features, and (2) blindly rank the top 3 features the model used to make predictions without accessing LIME explanations. Finally, LIME explanations were revealed and physicians were asked to rate their satisfaction on a 5-stars Likert scale. Results indicated that physicians agreed with model predictions (87%), and that LIME explanations were equal to the physicians’ explanations in at least two features for all patients (69%). Also, physicians were satisfied with LIME predictions (78%), and 68% presented a positive attitude towards explainability in terms of trust. However, such positive attitudes should not be translated into increased trust, as they found inconsistency in LIME explanations. Similarly in the scope of sepsis prediction, Sivaraman et al. (2023) examined clinicians’ interactions with an AI-based interpretable treatment recommender. The system was modeled using XGBoost, and SHAP as feature explainer. Through a think-aloud study involving 24 intensive care clinicians, the analysis revealed that explanations improved clinicians’ perceptions of the AI’s usefulness. In particular, the study identified four distinct behavior patterns among clinicians summarised as: Ignore, Negotiate, Consider, and Rely. Overall, clinicians who found the AI recommendations most useful saw it as additional evidence alongside their assessments.
In the field of medical wearable technologies, Kim et al. (2022) developed MindScope, a system based on a personalized ML system trained on an XGBoost to predict patients’ stress levels. Interviews were conducted to gather feedback on the relevancy of feature importance (SHAP values) in three different explanation settings: (1) no explanations, (2) broad explanations, and (3) detailed explanations. Overall, it was found that the explanation-based settings were more useful than no explanations. However, detailed explanations were perceived as equally useful as broad explanations, thus questioning how detailed explanations should be.
Next, Matthiesen et al. (2021) conducted a human study in the application of a CDSS able to predict ventricular fibrillation. The study was organized into three phases: (1) observations were conducted at the clinic to comprehend the clinical workflow, (2) development of the CDSS adopting a random forest model to predict the risk score, showing the top/bottom 5 most important features through LIME, and (3) interview study to assess feasibility. Although many clinicians did not change their decisions, they claimed that having a prediction tool could increase confidence in their actions. Also, the visualization of key features was useful, as the presentation of key features facilitates explainability regarding factors contributing to the risk of disease. However, the explainability component did not alter clinical actions.
In endocrinology, Kovalchuk et al. (2022) proposed a three-staged CDSS to for diabetic diagnosis as follows: (1) Basic Reasoning, i.e., implementing healthcare practices by referencing existing policies to guide decision-making, (2) Data-Driven Predictive Modeling, i.e., applying machine learning models in the context of the previous stage, and (3) Explanation, in which XAI techniques are applied to provide better recommendations. The authors tested the framework in a diabetes mellitus case study. In the first stage, expert endocrinologists guided the selection of insulin resistance predictors. In the second stage, a predictive model was optimized to estimate the risk. Finally, SHAP values were employed to interpret the CDSS predictions, which were then validated using questionnaires. Unexpectedly, though not extensively, the basic reasoning case showed higher understandability, followed by the explanation case, and the data-driven case. These findings suggest that adding explanations is useful, but well-grounded, known conventional clinical practices may still outperform AI-driven solutions.
Finally, integrating large language models (LLMs), Rajashekar et al. (2024) developed GutGPT, a system tailored to provide a natural language interface for gastrointestinal bleeding prediction, including hospital-based intervention. This system integrates LLMs interface, with a standard random forest model to predict a risk score. Then, interpretability plots such as individual conditional expectations (ICE) and accumulated local effects (ALE) were used to explain feature-wise effects on predicted scores. By interacting with human text, GutGPT leverages in-context learning to present easy-to-understand natural language explanations alongside XAI explanations for predictions. System usability was measured by means of interviews and surveys post hoc, finding that perceptions were mostly positive.
3.3 CDSS adopting XAI Methodologies in Deep Learning
CDSS using deep learning mainly rely on computer vision architectures to locally explain predictions, usually in form of heatmaps. Based on the explainability methodology, dependency may be either agnostic or specific. For example, agnostic methodologies such as integrated gradients could be applied to wider families of deep learning models; while, specific methodologies may be reserved for only a family of models (e.g., GradCAM for CNNs only).
In the field of model-agnostic XAI methodologies, Ellenrieder et al. (2023) extended the work of Pumplun et al. (2023) that leveraged a DenseNet and the integrated gradients framework to produce heatmaps of image regions to highlight cancer malignancy probabilities. Working with radiologists, a think-aloud study was conducted to assess how much they would learn with four differently developed UIs: non-explainable high-performing (NEHP), explainable high-performing (EHP), non-explainable low-performing (NELP), and explainable low-performing (ELP). Four hypotheses were confirmed: H1 - small learning gains occur with NEHP; H2 - larger learning gains occur with EHP; H3 - false learning happens with NELP; H4 - false learning is reduced with ELP, suggesting that explanations can significantly enhance clinical decision-making through learning.
Next, model-specific XAI methodologies were presented in the works from Hwang et al. (2022), Sabol et al. (2020), Singh et al. (2021), and Neves et al. (2021). Here, Hwang et al. (2022) developed a CDSS to aid sleep-staging technicians in reviewing AI-generated sleep staging results. The authors developed the study in three phases. Firstly, sleep technicians were asked why they would need explanations. Secondly, a user observation study was performed to understand the sleep staging conventions of clinical practitioners. Thirdly, a CDSS based on CNNs was developed, with convolutional filters, saliency values, and intermediate activation, as sources of information for generating explanations. Human evaluations suggested that explanations from the system effectively served as convincing elements for model predictions; however, they failed to reveal ambiguous predictions. As a general principle, technicians sought explanations to confirm the accuracy of the AI predictions, while adhering to their clinical knowledge.
Next, in the field of oncology, Sabol et al. (2020) adopted a Cumulative Fuzzy Class Membership Criterion (CFCMC) to predict colorectal cancer. They developed two segmentation systems for whole-slide histopathological images: one using a standalone CNN and the other using the CFCMC classifier on CNN-extracted features to provide explanations. Through surveys with 14 pathologists, the CFMC showed higher acceptability and trust compared to a plain CNN, pointing out that explainable classifiers improve AI usability and reliability in medical settings.
Next, Singh et al. (2021) evaluated multiple types of different visual explanation methods for an Inception-v-3 model trained to diagnose eye retinal diseases, while Neves et al. (2021) similarly adopted various explainability methods to analyze how a CDSS based on electrocardiogram (ECG) time series data could be used to predict heart arrhythmia, pointing out which explanatory method would perform optimally. In both studies, practitioners were carefully surveyed, resulting in an attribution method based on Taylor series expansions (DeepTaylor) and LIME, respectively, to be considered useful, in particular by less experienced clinicians. Finally, achieving comparable usability outcomes through semi-structured interviews, Zhang et al. (2024) uniquely leveraged an attention module to generate a fixed-length vector fed as input to an LSTM to select important variables, with attention weights to be used as the feature’s importance score, to develop SepsisLab, a CDSS able to predict ahead-of-time sepsis. Clinicians suggested that providing more explanations could improve their decision-making, but shifting the AI focus away from final decision predictions could improve their overall experience.
4 Challenges and Future Directions
4.1 How XAI Impacts CDSS: a Socio-Technical Perspective
We have reviewed the diversity of CDSS that adopt XAI human-centered evaluations in the medical field. Given the high-risk environment, application-grounded evaluations with domain experts were the main evaluation methodology. On average, clinical perceptions of XAI-based CDSS were mostly positive. For example, Ellenrieder et al. (2023) demonstrated that providing explanations not only enhances clinicians’ CDSS learning but also reduces false learning. In particular, Neves et al. (2021) demonstrated that less experienced practitioners could benefit from XAI explanations, while Brennan et al. (2019) revealed that adding XAI aids in significant improvements in physicians’ risk assessments.
However, incorporating XAI still poses significant challenges. For instance, Barda et al. (2020) discovered that the capacity to absorb explanations varied across clinical roles and levels of AI knowledge, thus advocating solutions that demand less cognitive effort. For the same reason, Kim et al. (2022) found that detailed explanations were perceived as equally useful as broad explanations, thus questioning how deep explanations should extend. Again, Kyrimi et al. (2020) documented that too-engineered explanations led to brain clutter, preventing their full internalization.
Notwithstanding, Sivaraman et al. (2023) more granularly identified that clinicians who ignored explanations were already confident in their own settled decisions, but clinicians who “negotiated” accepting at least one aspect of the recommendation were able to reach a balanced intermediate solution that combined with their own intuition. Similarly, Matthiesen et al. (2021) argued that ML and explainability did not surpass specialized and experienced clinicians’ capabilities, but at best support them during the clinical workflow, as well as Kyrimi et al. (2020), who found that, although XAI was useful, current conventional clinical practices are still perceived as more understandable. Finally, we highlight that only Hwang et al. (2022); Matthiesen et al. (2021) early gathered clinicians’ feedback before CDSS development, which may be crucial to ensure that the CDSS aligns with real-world clinical expectations, thus purportedly reducing frictions in subsequent stages. These findings emphasize the need for explanations to align with prior clinical knowledge through early stakeholder feedback and to minimize the cognitive burden on clinicians using CDSS.
More in general, despite explainability methods face intrinsical issues such as lack of robustness Slack et al. (2020) and intra-method disagreement Krishna et al. (2022), issues that may be framed as independent research topics per sé, the friction from clinicians in perceiving explainability may be attributed to socio-technical gaps between human requirements and the effective utility of technological solutions, because human activity is “flexible and nuanced”, while computational mechanisms are “rigid and brittle” Ackerman (2000). In other words, a socio-technical gap refers to the discrepancy between human needs, including expectations, and the technological capabilities designed to support them, highlighting the limitations of technological systems in fully accommodating the nuances of human interactions. This general theoretical construct can be extended to the purpose of XAI explanations in CDSS, as explanation capabilities might not meet rigorous human requirements from precise clinical knowledge, given that CDSS are dramatically at high stake, as output recommendations may have a significant impact on patient lives. In other words, implementing XAI solutions in CDSS might not always bridge the socio-technical gap when compared to systems that do not adopt XAI, or systems that use basic reasoning in current clinical knowledge.
Therefore, we suggest a reference protocol for effective CDSS development. This is an iterative process that requires exploring and connecting both the XAI and HCI fields, leveraging their theoretical constructs to eventually devise a practical methodology. We begin by identifying two theoretical dimensions: T1: determining stakeholders and their requirements for use case models (human requirement realism), and T2: developing XAI evaluation methods to reliably assess whether and to what extent these requirements are met in downstream use cases (context realism) Liao and Varshney (2021). Consistent with Doshi-Velez and Kim (2017), who stated that in high-stake domains, application-grounded XAI evaluations should be preferred because of their high fidelity, T2 particularly entails trade-offs with higher pragmatic and monetary costs to bear for effective evaluation.
In practical terms, designing a human-centered CDSS that conforms to the above-mentioned principles involves creating a system that is accurate in treatment recommendations, explainable in how and why such recommendations are conveyed, easily usable by clinicians, thus not imposing a significant cognitive burden, and seamlessly integrable into extant technological workflows. In other words, all stakeholders’ requirements (human requirement realism) should be addressed thoroughly by the CDSS use case (context realism).
4.2 A Medical Field-Agnostic, Stakeholder-Centric Iterative Approach to CDSS Development
The iterative process can be decomposed into the following phases: (1) stakeholder identification, (2) stakeholders’ requirements identification, and (3) system development and evaluation.
First, it is fundamental to identify all relevant stakeholders in each specific case. As a rule of thumb, identifiable stakeholders in the healthcare domain may be: (1) patients, (2) clinicians, (3) clinics and hospitals - including their corporate governance and management, and (4) CDSS developers - including researchers and engineers. Patients may be considered as passive stakeholders, as they do not directly interact, but are “targeted” by the system. On the contrary, clinicians are actively supported by CDSS, corporate management may decide whether the clinic/hospital should adopt AI-based solutions, evaluating the costs and benefits, and developers are responsible for the creation and maintenance of the technological workflows.
Second, we argue that a participatory AI methodology in which all stakeholders are early on involved is an effective starting point, where diverse requirements are early considered, potentially reducing frictions afterward, e.g., Hwang et al. (2022); Matthiesen et al. (2021). This entails gathering all stakeholders’ needs, setting goals, and clearly framing an evaluation strategy specific to the CDSS and to the medical field in which the CDSS would operate. To this end, we suggest developers inform and instruct non-technical stakeholders of how CDSS operate ex-ante, including carefully describing their inner workings based on machine learning model(s) and XAI explainability component(s), and possible evaluation frameworks. Here, presenting them with relevant taxonomies can help them become familiar with the material more effectively (see Section 2.2 for XAI principles and methods, and 2.3 for human-centered evaluations). Methodologies may include a pilot study for existing prototypes or a formative study if none exist. This phase should capture stakeholders’ expectations and assess the need for specialized evaluation methods in the CDSS’s medical field. Also, focus groups can be particularly useful, as they convene all stakeholders simultaneously.
Third, CDSS developers would optimally internalize stakeholders’ requirements gathered from the previous phase to develop a prototype to be quickly tested with synthetic patients’ data. The prototype should adhere to both technical backend requirements and usability requirements and be designed with a focus on future adoption, thus seamlessly integrable into medical workflows. We argue that concurrent think-aloud studies can be effective in this stage, as clinicians (the end users) can articulate their thoughts on the perception of the accuracy and explainability of the models.
In general, steps 2 and 3 should be conceived as iterative, adhering to the concept that continuous feedback stimulates CDSS improvement. This iterative process guarantees that system performance is continuously and rigorously improved. Through the incorporation of ongoing feedback, the CDSS can be fine-tuned with new information, thus refining both its predictive value and clinical usefulness. As such, this refinement process facilitates evidence-based decision-making and optimizes overall system effectiveness in clinical practice.
4.3 Conclusion
In this work, we have examined the role of explainability in CDSS and the socio-technical challenges it faces. While explainability enhances clinicians’ trust, its effectiveness varies based on user expertise, cognitive load, and the depth of explanation required. The socio-technical gap remains a central challenge, as human decision-making is inherently flexible, whereas (X)AI models operate within rigid constraints. Given the high stakes in healthcare, rigorous evaluation strategies - particularly application-grounded methodologies - are necessary to ensure that CDSS align with stakeholders’ requirements. To address these challenges, we proposed an iterative framework for CDSS development to integrate XAI and HCI principles. By identifying key stakeholders, gathering their requirements through participatory methodologies, and continuously refining CDSS through iterative evaluations, we advocate for a development approach that prioritizes usability, trust, and clinical integration. Overall, rather than replacing human expertise, CDSS should serve as augmentative tools that enhance clinicians’ decision-making, ultimately fostering more effective and informed patient care.
Acknowledgments
This work was funded by Fundação para a Ciência e a Tecnologia (UIDB/00124/2020, UIDP/00124/2020 and Social Sciences DataLab - PINFRA/22209/2016), POR Lisboa and POR Norte (Social Sciences DataLab, PINFRA/22209/2016).